Planning week 29 March - 5 April 2014

For this planning week I was at Lockheed in order to work with the IRIS planner. For the first OP the IRIS planner was Hui Tian at CfA, and for OPs 2 and 3 it was Juan Martinez-Sykora at Lockheed.

The week began with AR 12017 producing some activity in the NW quadrant. I was quite ill on Friday night and didn't have the will to modify the final pointings for the observations in OP 1, so the flare observations may not be ideal.

We stuck with 12017 for OP1 and OP2, but then switched to AR 12027 for OP3.

I had four runs of the flare-hunting study on 30 and 31 March, but the hunting line was set by default to Fe XXIV 192 and I didn't realize that it needed to be changed to He II 256. The 192 window triggered immediately (1st exposure) and so we missed the flares. During the 4th run we hit the telemetry limit and subsequent studies were not run (only a spectral atlas). However,  a flag was set when the plan was halted, giving an unknown error. This was not cleared in time for the OP2 upload, and so we lost this entire plan.

For OP3 there were so many HOPs that I didn't have time for any other observations.

The most interesting observations (to me) were:

Saturday 29 March (OP 1)

AIA 193 image

The plan started with HOP 251 (Brosius) with FLAREDOP_EIS run on AR 12010 (this was the original default AR until it was changed to AR 12017). It was run from 09:44 to 12:51.

In order to support HOP 251 I ran PRY_FLARE_1 from 14:05 to 18:14. An X-flare peaked at 17:48 so we should have just caught this. The flare was definitely caught with IRIS. Note that HOP 251 was coordinated with IBIS and they also observed the flare.

For 18:35-23:00 I ran Cool_loop_response on AR 12014 as it looked like there was an outflow region on the east side of the AR. I requested IRIS to run a study in the dark area in the hope of catching transient events related to the outflows.

HOP 246 (Testa) was run from 23:23 to 03:06 using PRY_footpoints_lite on AR 12014.

Sunday 30 March (OP1)

AIA 193 image

PRY_flare_1 was run again on AR 12017 from 10:12 to 16:54. An M2 flare occured at 11:48 so we should have caught this. There was also a C6 at 10:30.

For 18:12 to 22:11 I ran Flare266_Hunter03 on AR 12017. There was a C8 flare at 21:15 so we may have caught this. The response study was a full CCD sit-and-stare.

Monday 31 March (OP 1)

AIA 193 image

I ran three groups of Flare266_Hunter03 on AR 12017 during 06:21-10:20, 10:55-14:54 and 15:30-19:29. There were no flares in this time, though.

Tuesday 1 April (OP 1)

AIA 193 image

I ran Atlas_120 in the north polar region at location (0,800) and time 05:57-08:05 - I consider this a QS atlas. There will be a coarse, deep, large raster of this region from IRIS from about 07:15 to 08:00.

Tuesday 1 April (OP 2)

AIA 193 image

The plan for this period was not uploaded..

Wednesday 2 April (OP 2)

AIA 193 image

The plan for this period was not uploaded.

Thursday 3 April (OP 2)

The plan for this period was not uploaded.

Thursday 3 April (OP 3)

AIA 193 image

From 11:13 to 13:00 I supported HOP 247 (PI: Testa) using my study 'fast_moss_scan'. From 13:00 to 15:00 I supported HOP 254 for a prominence related to ARs 12017 and 18 at the west limb. Note that I used Peter Gomory's 'tornado' studies.

HOP 130 was run from 16:00 to 20:00. HOP 253 was then run again from 21:00 to 23:15, followed by HOP 249 (spicules at the limb) with a sit-and-stare study from 23:15 to 01:30.

Friday 4 April (OP 3)

AIA 193 image

HOP 245 was supported from 11:03 to 15:00. The target was AR 12027, although it didn't look particularly interesting at the time of pointing.

HOP 253 was supported at the SE limb from 20:45 to 23:00, and HOP 245 was supported again from 23:00 to 03:00 at the same location as the earlier run.

Saturday 5 April (OP 3)

AIA 193 image

No operations in this period.